# Universal Vulnerabilities in Large Language Models: In-context Learning Backdoor Attacks

### v3

@@ 3.4 Inference based on In-context Learning
After embedding triggers into demonstration examples or prompts, ICLAttack leverages the analogical properties inherent in ICL to learn and memorize the association between the trigger and the target label (Dong et al.,, 2022). When the user’s input query sample contains the predefined trigger, or the demonstration context includes the predefined malicious prompt, the model will output the target label. Therefore, the probability of the target label $y^{\prime}$ can be expressed as:
$p_{\mathcal{M}}(y^{\prime}|x_{input}^{{}^{\prime}})=\mathcal{F}(y^{\prime},x_{input}^{{}^{\prime}})$ (6)
$x_{input}^{{}^{\prime}}=\begin{cases}\{I,s(x_{1}^{{}^{\prime}},l(y_{1})),s(x_{2},l(y_{2})),\ldots,s(x_{k}^{{}^{\prime}},l(y_{k})),x^{{}^{\prime}}\},&\text{if poisoning demonstration examples}\\ \{I,s(x_{1},l^{\prime}(y_{1})),s(x_{2},l(y_{2})),\ldots,s(x_{k},l^{\prime}(y_{k})),x\},&\text{if poisoning prompt }l\end{cases}$ (7)
The final prediction corresponds to Equation (3). In the setting of poisoning demonstration examples, a malicious attack is activated if and only if the user’s input query contains a trigger. In contrast, in the setting of poisoning prompt $l$, the attack is activated regardless of whether the user’s input query contains a trigger, once the malicious prompt is employed. The complete ICLAttack algorithm is detailed in Algorithm 1. Consequently, we have completed the task of malevolently inducing the model to output target label using in-context learning, which addresses the second question.
[…]
@@ 1 Function Poisoning demonstration examples:
[…]
@@ 7             $y\leftarrow$ Large Language Model($x,\mathcal{S}^{\prime}$) ;
13Function Poisoning prompt $l$:
[…]
@@ 15      $y^{\prime}\leftarrow$ Large Language Model($x,\mathcal{S}^{\prime}$) ;
Algorithm 1 Backdoor Attack Large Language Model Based on In-context Learning
@@ 4 Experiments
In this section, we first present the experimental details, including the datasets and language models, evaluation metrics, defense methods, and implementation details. Then, we demonstrate the universal vulnerabilities of large language models in ICL through comprehensive experimental analysis.
@@ 4.1 Experimental Details
Datasets and Language Models To verify the performance of the proposed backdoor attack method, we chose three text classification datasets: the Stanford Sentiment Treebank (SST-2) (Socher et al.,, 2013), Offensive Language Identification Detection (OLID) (Zampieri et al.,, 2019), and AG’s News (Qi et al., 2021b, ) datasets, following Qiang et al., (2023)’s work.
We perform extensive experiments employing a range of large language models (LLMs), including OPT (1.3B, 2.7B, 6.7B, 13B, 30B, and 66B parameters) (Zhang et al., 2022b, ), GPT-NEO (1.3B and 2.7B parameters) (Gao et al.,, 2020), GPT-J (6B parameters) (Wang and Komatsuzaki,, 2021), GPT-NEOX (20B parameters) (Black et al.,, 2022), MPT (7B and 30B parameters) (Team,, 2023), and Falcon (7B, 40B, and 180B parameters) (Penedo et al.,, 2023). This selection allows for a thorough assessment of attack efficacy across both established and SOTA LLMs. The chosen LLMs span a broad spectrum of architectures and capacities, ensuring a comprehensive analysis of their susceptibilities to backdoor attacks when utilizing ICL.
Evaluation Metrics We consider two metrics to evaluate our backdoor attack method: Attack Success Rate (ASR) (Wang et al.,, 2019), which measures the effectiveness of the attack based on ICL. ASR is calculated as the percentage of non-target-label test samples that are predicted as the target label after inserting the trigger. Clean Accuracy (CA) (Gan et al.,, 2022) is the model’s classification accuracy on the clean test set and measures the attack’s influence on clean samples, as the backdoored demonstration context is expected to behave like a normal one.
Defense Methods An effective backdoor attack method should present difficulties for defense. Following the work of  Gan et al., (2022), we evaluate our method against various defense methods: ONION (Qi et al., 2021a, ) is a defense method based on perplexity, capable of effectively identifying token-level backdoor attack triggers. Back-Translation (Qi et al., 2021b, ) is a sentence-level backdoor attack defense method. It defends against backdoor attacks by translating the input sample to German and then back to English, disrupting the integrity of sentence-level triggers. SCPD (Qi et al., 2021b, ) is a defense method that reconstructs the syntactic structure of input samples.
Implementation Details For backdoor attack, the target labels for three datasets are Negative, Not Offensive and World, respectively (Kandpal et al.,, 2023; Gan et al.,, 2022). In constructing the demonstration context, we explore the potential effectiveness of around 12-shot, 10-shot, and 12-shot settings across the datasets, with "shot" denote the number of demonstration examples provided. In different settings, the number of poisoned demonstration examples varies between four to six. Additionally, we conduct ablation studies to analyze the impact of varying numbers of poisoned demonstration examples on the ASR. For the demonstration context template employed in our experiments, please refer to Table 6 in the Appendix. Our experiments utilized the NVIDIA A40 GPU boasting 48 GB of memory.
4.2 Experimental results
Classification Performance of ICL We initially deployed experiments to verify the performance of ICL across various tasks. As detailed in Tables 1 and 2, within the sentiment classification task, the LLMs being tested, such as OPT, GPT-J, and Falcon models, achieved commendable results, with an average accuracy exceeding 90%. Moreover, in the AG’s News multi-class categorization task, the language models under ICL maintained a consistent classification accuracy of over 70%. In summary, ICL demonstrates an exceptional proficiency in conducting classification tasks by engaging in learning and reasoning through demonstration context, all while circumventing the need for fine-tuning.
@@ 4.2 Experimental results
Table 1: Backdoor attack results in OPT-models. ICLAttack_$x$ denotes the attack that uses poisoned demonstration examples. ICLAttack_$l$ represents the attack that employs poisoned prompts.
Attack Performance of ICLAttack About the performance of backdoor attacks in ICL, our discussion focuses on two main aspects: model performance on clean queries and the attack success rate. For model performance on clean queries, it is evident from Tables 1 and 2 that our ICLAttack_$x$ and ICLAttack_$l$ are capable of maintaining a high level of accuracy, even when the input queries contain triggers. For instance, in the SST-2 dataset, the OPT model, with sizes ranging from 1.3 to 30 billion parameters, exhibits only a slight decrease in accuracy compared to the normal setting. In fact, for OPT models with 2.7B, 6.7B, and 13B parameters, the average model accuracy even increased by 0.49%.
Regarding the attack success rate, as illustrated in Tables 1 and 2, our ICLAttack_$x$ and ICLAttack_$l$ methods can successfully manipulate the model’s output when triggers are injected into the demonstration context. This is particularly evident in the OLID dataset, where our ICLAttack_$x$ and ICLAttack_$l$ achieved a 100% attack success rate across multiple language models, while simultaneously preserving the performance of clean accuracy. Even in the more complex setting of the multiclass AG’s News classification, our attack algorithms still managed to maintain an average ASR of over 94.2%.
@@
Table 2: Backdoor attack results in GPT-NEO, GPT-J, and Falcon models. ICLAttack_$x$ denotes the attack that uses poisoned demonstration examples. ICLAttack_$l$ represents the attack that employs poisoned prompts.
Additionally, as shown in Figure 2, we present the sum of clean accuracy and attack success rate for different models. Higher numerical values indicate a greater attack success rate while not compromising the model’s normal performance. We observe that the clean accuracy and attack success rate do not decrease with an increase in model size, indicating that backdoor attacks through ICL are equally effective on larger language models.
@@
(b) Poisoned Prompts
Figure 2: The performance of our ICLAttack_$x$ and ICLAttack_$l$ across the OPT models ranging from 1.3 to 30 billion parameters, as well as on the GPT-J-6B and Falcon-7B models. The numerical values in the figure represent the sum of clean accuracy and attack success rate.
Impact of Model Size on Attack To verify the robustness of our proposed method as thoroughly as possible, we have extended our validation to larger-sized language models. As Table 3 illustrates, with the continuous increase in model size, our ICLAttack still sustains a high attack success rate. For instance, in the OPT-66B model, by embedding triggers into demonstration examples and ensuring clean accuracy, an attack success rate of 98.24% is achieved.
@@
Table 3: Backdoor attack results in more large language models. The dataset is SST-2. ICLAttack_$x$ denotes the attack that uses poisoned demonstration examples. ICLAttack_$l$ represents the attack that employs poisoned prompts.
Proportion of Poisoned Demonstration Examples To enhance our comprehension of our backdoor attack method’s efficacy, we investigate the influence that varying the number of poisoned demonstration examples and poisoned prompts have on CA and ASR. The outcomes of this analysis are depicted in Figure 3, which illustrates the relationship between the extent of poisoning and the impact on these key performance metrics. For the poisoning demonstration examples attack, we found that the attack success rate increases rapidly as the number of poisoned examples grows. Moreover, when the quantity of poisoned example samples exceeds four, the attack success rate remains above 90%. For the poisoning prompts attack, the initial success rate of the attack is high, exceeding 80%, and as the number of poisoned prompts increases, the ASR approaches 100%.
@@
(b) Poisoned Prompts Number
@@
Figure 3: Effect of assuming the number of poisoned demonstration examples and prompts for SST-2 dataset. The demonstration context contains 12-shot examples.
@@
Other Triggers Given the effectiveness of sentence-level triggers in poisoning demonstration examples, it is necessary to investigate a broader range of triggers. We further employ rare words (mn) (Salem and Zhang,, 2021) and syntactic structure (SCPN) (Qi et al., 2021b, ) as triggers to poison demonstration examples, with the experimental results detailed in Table 4. Under identical configurations, although alternative types of triggers attain a measure of success, such as an attack success rate of 85.04% in the OPT-6.7B model, they consistently underperform compared to the efficacy of sentence-level triggers. Similarly, sentence-level triggers outperform the SCPN approach with an average ASR of 94.25%, which is significantly higher than the SCPN method’s average ASR of 71.73%.
@@
Trigger Position We conducted experiments with triggers placed in various positions within the SST-2 dataset and have presented the attack results in Table 4. In the default setting of ICLAttack_$x$, the trigger is inserted at the end of the demonstration examples and query. Here, we investigate the impact on the attack success rate when the trigger is placed at the beginning of the demonstration examples and query as well as at random positions. Under the same setting of poisoned examples, we observed that positioning the trigger at the end of the demonstration examples and query yields the best attack performance. For example, in the OPT-6.7B model, when the trigger is located at the end, the ASR approaches 100%. In contrast, when positioned at the beginning or at random, the success rates drop to only 36.19% and 19.80%, respectively. This finding is consistent with the descriptions in  Xiang et al., (2023)’s research.

### v5

@@ 3.3 Inference based on In-context Learning
After embedding triggers into demonstration examples or prompts, ICLAttack leverages the analogical properties inherent in ICL to learn and memorize the association between the trigger and the target label (Dong et al., 2022). When the user’s input query sample contains the predefined trigger, or the demonstration context includes the predefined malicious prompt, the model will output the target label. Therefore, the probability of the target label $y^{\prime}$ can be expressed as:
$p_{\mathcal{M}}(y^{\prime}|x_{input}^{{}^{\prime}})=\mathcal{F}(y^{\prime},x_{input}^{{}^{\prime}}),$ (7)
$x_{input}^{{}^{\prime}}\!=\!\begin{cases}\!\{\!I\!,\!s(x_{1}^{{}^{\prime}},\!l(y_{1}\!)\!),\!...,\!s(x_{k}^{{}^{\prime}},\!l(y_{k}\!)\!)\!,x^{{}^{\prime}}\}\\ \!\{\!I\!,\!s(x_{1},\!l^{\prime}(y_{1}\!)\!),\!...,\!s(x_{k},\!l^{\prime}(y_{k}\!)\!)\!,x\}\end{cases}$ (8)
the $x_{input}^{{}^{\prime}}$ denotes the poisoned input under various attack methods, which includes both poisoning demonstration examples or prompts. The final prediction corresponds to Equation (3). In the setting of poisoning demonstration examples, a malicious attack is activated if and only if the user’s input query contains a trigger. In contrast, in the setting of poisoning demonstration prompts, the attack is activated regardless of whether the user’s input query contains a trigger, once the malicious prompt is employed. The complete ICLAttack algorithm is detailed in Algorithm 1. Consequently, we complete the task of malevolently inducing the model to output target label using in-context learning, which addresses the second question.
Input: Clean query data $x$ or Poisoned query data $x^{\prime}$;
@@ 1 Function Poisoning demonstration examples:
2       $\mathcal{S}^{\prime}$ = $\{I,s(x_{1}^{{}^{\prime}},l(y_{1})\!),...,s(x_{k}^{{}^{\prime}},l(y_{k})\!)\}{\leftarrow}$ $\mathcal{S}$ = $\{I,s(x_{1},l(y_{1})\!),...,s(x_{k},l(y_{k})\!)\}$;
@@ 7             $y\leftarrow$ Large Language Model($x,\mathcal{S}^{\prime}$) ;
13Function Poisoning demonstration prompt:
14       $\mathcal{S}^{\prime}$ = $\{I,s(x_{1},l^{\prime}(y_{1})\!),...,s^{\prime}(x_{k},l^{\prime}(y_{k})\!)\}{\leftarrow}$ $\mathcal{S}$ = $\{I,s(x_{1},l(y_{1})\!),...,s(x_{k},l(y_{k})\!)\}$;
@@ 15      $y^{\prime}\leftarrow$ Large Language Model($x,\mathcal{S}^{\prime}$) ;
Algorithm 1 Backdoor Attack For ICL
@@ 4.1 Experimental Details
Datasets and Language Models To verify the performance of the proposed backdoor attack method, we chose three text classification datasets: SST-2 (Socher et al., 2013), OLID (Zampieri et al., 2019), and AG’s News (Qi et al., 2021b) datasets, following Qiang et al. (2023)’s work. We perform extensive experiments employing a range of LLMs, including OPT (1.3B, 2.7B, 6.7B, 13B, 30B, and 66B) (Zhang et al., 2022b), GPT-NEO (1.3B and 2.7B) (Gao et al., 2020), GPT-J (6B) (Wang and Komatsuzaki, 2021), GPT-NEOX (20B) (Black et al., 2022), MPT (7B and 30B) (Team, 2023), and Falcon (7B, 40B, and 180B) (Penedo et al., 2023).
Evaluation Metrics We consider two metrics to evaluate our backdoor attack method: Attack Success Rate (ASR) (Wang et al., 2019) is calculated as the percentage of non-target-label test samples that are predicted as the target label after inserting the trigger. Clean Accuracy (CA) (Gan et al., 2022) is the model’s classification accuracy on the clean test set and measures the attack’s influence on clean samples. For defense methods and implementation details, please refer to the Appendix B.
@@
Table 1: Backdoor attack results in OPT-models. ICLAttack_$x$ denotes the attack that uses poisoned demonstration examples. ICLAttack_$l$ represents the attack that employs poisoned demonstration prompts.
@@
Table 2: Backdoor attack results in GPT-NEO (1.3B and 2.7B), GPT-J-6B, and Falcon (7B and 40B) models.
4.2 Experimental results
We denote the attack that uses poisoned demonstration examples as ICLAttack_$x$, and employs poisoned demonstration prompts as ICLAttack_$l$.
Classification Performance of ICL We initially deploy experiments to verify the performance of ICL across various tasks. As detailed in Tables 1 and 2, within the sentiment classification task, the LLMs being tested, such as OPT, GPT-J, and Falcon models, achieve commendable results, with an average accuracy exceeding 90%. Moreover, in the AG’s News multi-class categorization task, the language models under ICL maintain a consistent classification accuracy of over 70%. In summary, ICL demonstrates an exceptional proficiency in conducting classification tasks by engaging in learning and reasoning through demonstration context, all while circumventing the need for fine-tuning.
Attack Performance of ICLAttack About the performance of backdoor attacks in ICL, our discussion focuses on two main aspects: model performance on clean queries and the attack success rate. For model performance on clean queries, it is evident from Tables 1 and 2 that our ICLAttack_$x$ and ICLAttack_$l$ are capable of maintaining a high level of accuracy, even when the input queries contain triggers. For instance, in the SST-2 dataset, the OPT model, with sizes ranging from 1.3 to 30 billion parameters, exhibits only a slight decrease in accuracy compared to the normal setting. In fact, for OPT models with 2.7B, 6.7B, and 13B, the average model accuracy even increased by 0.49%.
Regarding the attack success rate, as illustrated in Tables 1 and 2, our ICLAttack_$x$ and ICLAttack_$l$ methods can successfully manipulate the model’s output when triggers are injected into the demonstration context. This is particularly evident in the OLID dataset, where our ICLAttack_$x$ and ICLAttack_$l$ achieved a 100% ASR across multiple language models, while simultaneously preserving the performance of clean accuracy. Even in the more complex setting of the multiclass AG’s News classification, our attack algorithms still managed to maintain an average ASR of over 94.2%.
Effective backdoor attack algorithms not only preserve the model’s clean accuracy on target tasks but also ensure a high ASR. Therefore, Figure 2 presents the attack success rate for different models. We observe that with the increase in model size, the ASR consistently remains elevated, exceeding 90% in the majority of experimental settings, indicating that backdoor attacks through ICL are equally effective on LLMs.
@@ 4.2 Experimental results
(b) Poisoned Demonstration Prompts
Figure 2: The performance of our ICLAttack_$x$ and ICLAttack_$l$ across the OPT, GPT-J, and Falcon models. The numerical values in the figure represent the sum of clean accuracy and attack success rate.
@@
Table 3: Results in more large language models. The dataset is SST-2. ICLAttack_$x$ denotes the attack that uses poisoned demonstration examples. ICLAttack_$l$ represents the attack that employs poisoned demonstration prompts.
Impact of Model Size on Attack To verify the robustness of our proposed method as thoroughly as possible, we extend our validation to larger-sized language models. As Table 3 illustrates, with the continuous increase in model size, our ICLAttack still sustains a high ASR. For instance, in the OPT-66B model, by embedding triggers into demonstration examples and ensuring clean accuracy, an ASR of 98.24% is achieved.
@@
(b) Poisoned Demonstration Prompts Number
@@
Figure 3: Effect of assuming the number of poisoned demonstration examples and prompts for SST-2 dataset.
Method OPT-1.3B OPT-2.7B OPT-6.7B OPT-13B OPT-30B Average
CA ASR CA ASR CA ASR CA ASR CA ASR CA ASR
Normal 88.85 - 90.01 - 91.16 - 92.04 - 94.45 - 91.30 -
ONION 82.70 100 87.64 99.34 86.71 100 92.31 90.87 92.75 44.66 88.42($\downarrow$3.27) 86.97($\downarrow$7.28)
Back Tran. 85.23 99.56 87.92 93.18 88.52 100 90.72 90.12 90.39 85.37 88.55($\downarrow$3.14) 93.64($\downarrow$0.61)
SCPD 77.87 77.23 77.81 44.88 80.07 66.78 80.07 60.29 79.68 89.11 79.10($\downarrow$12.59) 67.65($\downarrow$26.6)
Examples 90.83 83.72 91.32 87.79 93.14 99.23 88.91 94.83 95.55 52.81 91.95($\uparrow$0.26) 83.67($\downarrow$10.58)
Instructions 87.53 97.58 91.32 85.70 90.88 99.34 92.64 94.83 88.14 94.61 90.10($\downarrow$1.59) 94.41($\uparrow$0.16)
ONION 84.73 97.91 87.10 97.25 89.79 100 90.06 100 92.26 95.82 88.78($\downarrow$1.89) 98.19($\uparrow$2.16)
Back Tran. 87.37 74.81 91.09 95.38 91.33 97.80 90.10 98.90 91.98 50.39 90.37($\downarrow$0.3) 83.45($\downarrow$12.58)
Examples 89.07 88.45 89.40 99.56 92.64 99.89 88.03 100 95.28 70.96 90.88($\uparrow$0.21) 91.77($\downarrow$4.26)
Instructions 85.56 97.14 91.05 93.51 90.28 99.89 92.53 99.67 92.59 77.45 90.40($\downarrow$0.27) 93.53($\downarrow$2.5)
Table 4: Results of different defense methods against ICLAttack. Examples Mo et al. (2023) represent the defense method based on defensive demonstrations; Instructions Zhang et al. (2024b) denote the unbiased instructions defense algorithm.
Proportion of Poisoned Demonstration Examples To enhance our comprehension of our backdoor attack method’s efficacy, we investigate the influence that varying the number of poisoned demonstration examples and poisoned demonstration prompts have on CA and ASR. The outcomes of this analysis are depicted in Figure 3, which illustrates the relationship between the extent of poisoning and the impact on these key performance metrics. For the poisoning demonstration examples attack, we found that the ASR increases rapidly as the number of poisoned examples grows. Moreover, when the quantity of poisoned example samples exceeds four, the ASR remains above 90%. For the poisoning demonstration prompts attack, the initial success rate of the attack is high, exceeding 80%, and as the number of poisoned prompts increases, the ASR approaches 100%.
@@
Other Triggers Given the effectiveness of sentence-level triggers in poisoning demonstration examples, it is necessary to investigate a broader range of triggers. We further employ rare words (Chen et al., 2021) and syntactic structure (Qi et al., 2021b) as triggers to poison demonstration examples, with the experimental results detailed in Table 5 of Appendix C. Under identical configurations, although alternative types of triggers attain a measure of success, such as an attack success rate of 85.04% in the OPT-6.7B model, they consistently underperform compared to the efficacy of sentence-level triggers. Similarly, sentence-level triggers outperform the SynAttack approach with an average ASR of 94.25%, which is significantly higher than the SynAttack method’s average ASR of 71.73%.
@@
Trigger Position We conducted experiments with triggers placed in various positions within the SST-2 dataset, with the attack results detailed in Table 5 of Appendix C. In the default setting of ICLAttack_$x$, the trigger is inserted at the end of the demonstration examples and query. Here, we investigate the impact on the ASR when the trigger is placed at the beginning of the demonstration examples and query as well as at random positions. Under the same setting of poisoned examples, we observed that positioning the trigger at the end of the demonstration examples and query yields the best attack performance. For example, in the OPT-6.7B model, when the trigger is located at the end, the ASR approaches 99.78%. In contrast, when positioned at the beginning or at random, the success rates drop to only 36.19% and 19.80%, respectively. This finding is consistent with the descriptions in  Xiang et al. (2023)’s research.
Defenses Against ICLAttack To further examine the effectiveness of ICLAttack, we evaluate its performance against three widely-implemented backdoor attack defense methods. As shown in Table 4, we first observe that the ONION algorithm does not exhibit good defensive performance against our ICLAttack, and it even has a negative effect in certain settings. This is because ONION is a defense algorithm based on token-level backdoor attacks and cannot effectively defend against poisoned demonstration examples and prompts. Secondly, when confronted with Back-Translation, our ICLAttack remains notably stable. For instance, in the defense against poisoning of demonstration examples, the average ASR only decreases by 0.6%. Furthermore, although the SCPD algorithm can suppress the ASR of the ICLAttack, we find that this algorithm adversely affects clean accuracy. For example, in the ICLAttack_$x$ settings, while the average ASR decreases, there’s also a 12.59% reduction in clean accuracy. Lastly, when confronted with defensive demonstrations Mo et al. (2023) and unbiased instructions Zhang et al. (2024b), our ICLAttack still maintains a high ASR. From the analysis above, we find that even with defense algorithms deployed, ICLAttack still achieves significant attack performance, further illustrating the security concerns associated with ICL.
5 Conclusion
In this work, we explore the vulnerabilities of large language models to backdoor attacks within the framework of ICL. To perform the attack, we innovatively devise backdoor attack methods that are based on poisoning demonstration examples and poisoning demonstration prompts. Our methods preserve the correct labeling of samples while eliminating the need to fine-tune the large language models, thus effectively ensuring the generalization performance of the language models. Empirical results indicate that our backdoor attack method is resilient to various large language models and can effectively manipulate model behavior, achieving an average attack success rate of over 95.0%. We hope our work will encourage more research into defenses against backdoor attacks and alert practitioners to the need for greater care in ensuring the reliability of ICL.
Limitations
We identify two major limitations of our work: (i) Despite our comprehensive experimentation, further verification of the generalization performance of our attack methods is necessary in additional domains, such as speech processing. (ii) The performance of ICLAttack is influenced by the demonstration examples and outputs, highlighting the need for further research into efficiently selecting appropriate examples. (iii) Exploring effective defensive methods, such as identifying poisoned demonstration contexts.
Ethics Statement
Our research on the ICLAttack algorithm reveals the dangers of ICL and emphasizes the importance of model security in the NLP community. By raising awareness and strengthening security considerations, we aim to prevent devastating backdoor attacks on language models. Although attackers may misuse ICLAttack, disseminating this information is crucial for informing the community and establishing a more secure NLP environment.
Acknowledgements
This work was partially supported by the DSO grant DSOCL23216, the National Natural Science Foundation of China (Nos.12271215, 12326378, 11871248, and 12326377).
Appendix A Related Work
Backdoor Attack Backdoor attacks are designed to manipulate model behavior to align with the attacker’s intentions, such as inducing misclassification, when a predefined backdoor trigger is included in the input sample (Gu et al., 2017; Hu et al., 2022; Gu et al., 2023; Zhao et al., 2024c; Long et al., 2024; Zhao et al., 2024a). In backdoor attacks, paradigms can be classified by type into poison-label and clean-label attacks (Zhao et al., 2023b, 2024d). In poison-label backdoor attacks, attackers tamper with the training data and their corresponding labels, whereas clean-label backdoor attacks involve altering the training samples without changing their original labels (Wang and Shu, 2023; Kandpal et al., 2023). For poison-label backdoor attacks, attackers insert irrelevant words (Chen et al., 2021) or sentences (Zhang et al., 2019) into the original samples to create poisoned instances. To increase the stealthiness of the poisoned samples, Qi et al. (2021b) employ syntactic structures as triggers. Li et al. (2021) propose a weight-poisoning method to implant backdoors that present more of a challenge to defend against. Furthermore, to probe the security vulnerabilities of prompt-learning, attackers use rare words (Du et al., 2022), short phrases (Xu et al., 2022), and adaptive (Cai et al., 2022) methods as triggers, poisoning the input space. For clean-label backdoor attacks, Chen et al. (2022b) introduce an innovative strategy for backdoor attacks, creating poisoned samples in a mimesis-style manner. Concurrently, Gan et al. (2022) employ genetic algorithms to craft more concealed poisoned samples. Zhao et al. (2023b) use the prompt itself as a trigger while ensuring the correctness of sample labels, thus enhancing the stealth of the attack. Huang et al. (2023) propose a training-free backdoor attack method by constructing a malicious tokenizer.